### v2

@@ C.2 Results
For the baseline results, I set $N=1,000$ and $T=2$, and use a second-order polynomial spline basis with its knot at the median for all series estimation steps. I run $M=1,000$ simulations. For each $i$, I construct $W_{i}$ as the mean over time of each coordinate of $X_{it}$. Following the theory, I set $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{{}^{\prime}}$. The performance of $\widehat{\overline{\omega}}$, $\widehat{\overline{b}}^{K},$ and $\widehat{\overline{b}}^{L}$ is summarized in Table 3. For notational simplicity, I use $\sum_{m}$ for $\sum_{m=1}^{M}$ and $\sum_{i,t,m}$ for $\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{m=1}^{M}$.
Table 3: Performance of $\widehat{\overline{b}}$
Formula $\widehat{\overline{\omega}}$ $\widehat{\overline{b}}^{K}$ $\widehat{\overline{b}}^{L}$
Bias $M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right)/\left|\overline{b}^{\left(d\right)}\right|$ 2.81% 3.22% 3.24%
rMSE $\sqrt{M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right)^{2}}/\left|\overline{b}^{\left(d\right)}\right|$ 2.89% 3.65% 3.62%
MND $M^{-1}\sum_{m}\left|\widehat{\overline{b}}_{m}^{\left(d\right)}-\overline{b}^{\left(d\right)}\right|/\left|\overline{b}^{\left(d\right)}\right|$ 2.81% 3.26% 3.26%
@@
The first row of Table 3 reports the normalized bias of each coordinate of $\widehat{\overline{b}}$. The bias is reasonably small across all three coordinates, with a magnitude between 2.81% and 3.24% of the length of the corresponding coordinate of $\overline{b}$. The second row shows the normalized rMSE of each coordinate of $\widehat{\overline{b}}$. My method achieves low normalized rMSEs between 2.89% and 3.65% of the size of the corresponding coordinate $\overline{b}^{\left(d\right)}$ of $\overline{b}$. The last row presents the normalized MNDs of each coordinate of $\widehat{\overline{b}}$. Again, the method performs well with an MND between 2.81% and 3.26% of the size of the corresponding coordinate $\overline{b}^{\left(d\right)}$ of $\overline{b}$.
@@
Next, I investigate the performance of $\widehat{b}_{t,m}\left(x\right)$, which is obtained by regressing the estimated $\widehat{b}_{1t,m}\left(\widehat{V}_{it,m},W_{i,m}\right)$ on the second-degree polynomial spline basis of $X_{it,m}$ with its knot at the median for each $\left(t,m\right)$ combination. For the true $b_{t}\left(x\right)$, due to the complex dependence structure of $X_{it}$ on $\beta_{it}$, there is no analytical solution available. Therefore, for each $t$ I pool the observations across all $M$ simulations ($N\times M$ observations for each $t$) and approximate $b_{t}\left(x\right)$ by regressing the true $\beta_{it,m}$ on the same spline basis functions of $X_{it,m}$.
@@
Table 4 presents the results. Note that by definition the bias of $\widehat{b}_{t}\left(x\right)$ is the same as $\widehat{\overline{b}}$ so I omit it here. The normalized rMSE of $\widehat{b}_{t}\left(x\right)$ is bigger than that of $\widehat{\overline{b}}$, with a magnitude between 3.36% and 5.59% of the size of the corresponding $\overline{b}^{\left(d\right)}$. The normalized MND follows a similar pattern. The performance of $\widehat{b}_{t}\left(x\right)$ for estimating $b_{t}\left(x\right)$ is not as good as that of $\widehat{\overline{b}}$ for $\overline{b}$, which is expected because (i) $b_{t}\left(x\right)$ is a function rather than a finite-dimensional vector and (ii) there is approximation error in calculating the true $b_{t}\left(x\right)$ via simulations.
Table 4: Performance of $\widehat{b}_{t}\left(x\right)$
Formula $\widehat{\omega}_{t}\left(x\right)$ $\widehat{b}_{t}^{K}\left(x\right)$ $\widehat{b}_{t}^{L}\left(x\right)$
rMSE $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left(\widehat{b}_{t,m}^{\left(d\right)}\left(x_{it,m}\right)-b_{t}^{\left(d\right)}\left(x_{it,m}\right)\right)^{2}}/\left|\overline{b}^{\left(d\right)}\right|$ 3.36% 5.57% 5.59%
MND $\left(NTM\right)^{-1}\sum_{i,t,m}\left|\widehat{b}_{t,m}^{\left(d\right)}\left(x_{it,m}\right)-b_{t}^{\left(d\right)}\left(x_{it,m}\right)\right|/\left|\overline{b}^{\left(d\right)}\right|$ 2.87% 4.36% 4.37%
@@
To show how robust my method is in estimating $\overline{b}$ and $b_{t}\left(x\right)$, I conduct another set of exercises. I evaluate the performance of my estimator using rMSE defined as $\sqrt{M^{-1}\sum_{m}\left\|\widehat{\overline{b}}_{m}-\overline{b}\right\|^{2}}/\left\|\overline{b}\right\|$ for the whole vector of $\overline{b}$, and $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left\|\widehat{b}_{t,m}\left(x_{it,m}\right)-b_{t}\left(x_{it,m}\right)\right\|^{2}}/\left\|\overline{b}\right\|$ for the whole vector of $b_{t}\left(x\right)$. First, I vary $N$ and $T$, and present the results in Table 5. As expected, a larger $N$ is good for overall performance. However, the magnitude in the improvement of performance is mild, possibly caused by the fact that with more agents I need more data to control for the increasing degree of heterogeneity. On the other hand, I find that the method performs reasonably well even with a small sample size of $N=500$. Having a larger $T$ improves the performance of my method, which is again expected as I can exploit more information from repeated observations of the same individual to better control for the fixed effect $A_{i}$.
Table 5: Performance under Varying $N$ and $T$
@@
Second, I vary the degree of the spline basis functions used to construct the series estimators for all steps, and present the results in Table 6. The knot is placed at the median for all cases. I find that increasing the degree of basis functions from one to two improves estimation accuracy significantly. When I increase the degree of basis functions from two to three, the performance deteriorates. I attribute it to the fact that there are many more regressors in each step of regression when $d=3$, leading to a possible over-fitting problem. Based on this result, I use the second-degree splines in the empirical application. [^29] One may also use the AIC criterion to select the degree of basis functions.
Table 6: Performance under Varying Degree of Basis Functions
@@
[…]
Table 7: Performance with and without Ex Post Shocks
Add $\epsilon_{it}$ to $Y_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
No 2.99% 3.69%
Yes 3.64% 5.43%
Add $\upsilon_{it}$ to $\left(\beta_{it}^{K},\beta_{it}^{L}\right)$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
No 2.99% 3.69%
Yes 3.57% 5.19%

### v3

@@ C.2 Results
For the baseline results, I set $n=1,000$ and $T=2$, and use a second-degree polynomial spline basis with its knot at the median for all series estimation steps. I run $M=1,000$ simulations. For each $i$, I construct $W_{i}$ as the mean over time of each coordinate of $X_{it}$. Following the theory, I set $V_{it}\coloneqq\left(F_{\left.k_{it}\right|Z_{it},W_{i}}\left(\left.k_{it}\right|Z_{it},W_{i}\right),F_{\left.l_{it}\right|Z_{it},W_{i}}\left(\left.l_{it}\right|Z_{it},W_{i}\right)\right)^{\prime}$. The performance of $\widehat{\overline{\omega}}$, $\widehat{\overline{b}}_{K},$ and $\widehat{\overline{b}}_{L}$ is summarized in Table A.5. For notational simplicity, I use $\sum_{m}$ for $\sum_{m=1}^{M}$ and $\sum_{i,t,m}$ for $\sum_{i=1}^{n}\sum_{t=1}^{T}\sum_{m=1}^{M}$.
Table A.5: Performance of $\widehat{\overline{b}}$
Formula $\widehat{\overline{\omega}}$ $\widehat{\overline{b}}_{K}$ $\widehat{\overline{b}}_{L}$
Bias $M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{d}^{\left(m\right)}-\overline{b}_{d}\right)/\left|\overline{b}_{d}\right|$ 2.81% 3.22% 3.24%
rMSE $\sqrt{M^{-1}\sum_{m}\left(\widehat{\overline{b}}_{d}^{\left(m\right)}-\overline{b}_{d}\right)^{2}}/\left|\overline{b}_{d}\right|$ 2.89% 3.65% 3.62%
MND $M^{-1}\sum_{m}\left|\widehat{\overline{b}}_{d}^{\left(m\right)}-\overline{b}_{d}\right|/\left|\overline{b}_{d}\right|$ 2.81% 3.26% 3.26%
@@
The first row of Table A.5 reports the normalized bias of each coordinate of $\widehat{\overline{b}}$. The bias is reasonably small across all three coordinates, with a magnitude between 2.81% and 3.24% of the length of the corresponding coordinate of $\overline{b}$. The second row shows the normalized rMSE of each coordinate of $\widehat{\overline{b}}$. My method achieves low normalized rMSEs between 2.89% and 3.65% of the size of the corresponding coordinate $\overline{b}_{d}$ of $\overline{b}$. The last row presents the normalized MNDs of each coordinate of $\widehat{\overline{b}}$. Again, the method performs well with an MND between 2.81% and 3.26% of the size of the corresponding coordinate $\overline{b}_{d}$ of $\overline{b}$.
@@
Next, I investigate the performance of $\widehat{b}_{t}\left(x\right)$, which is obtained following the estimation procedure outlined in Subsection 4.1. For the true $b_{t}\left(x\right)$, due to the complex dependence structure of $X_{it}$ on $\beta_{it}$, there is no analytical solution available. Therefore, for each $t$ I pool the observations across all $M$ simulations ($n\times M$ observations for each $t$) and approximate $b_{t}\left(x\right)$ by regressing the true $\beta_{it}$ on the same spline basis functions of $X_{it}$.
@@
Table A.6 summarizes the results. Note that by definition the bias of $\widehat{b}_{t}\left(x\right)$ is the same as that of $\widehat{\overline{b}}$, so I omit it here. The normalized rMSE of $\widehat{b}_{t}\left(x\right)$ is bigger than that of $\widehat{\overline{b}}$, with a magnitude between 3.36% and 5.59% of the size of the corresponding $\overline{b}_{d}$. The normalized MND follows a similar pattern. The performance of $\widehat{b}_{t}\left(x\right)$ for estimating $b_{t}\left(x\right)$ is not as good as that of $\widehat{\overline{b}}$ for $\overline{b}$, which is expected because (i) $b_{t}\left(x\right)$ is a function rather than a finite-dimensional vector and (ii) there is approximation error in calculating the true $b_{t}\left(x\right)$ via simulations.
Table A.6: Performance of $\widehat{b}_{t}\left(x\right)$
Formula $\widehat{\omega}_{t}\left(x\right)$ $\widehat{b}_{t,K}\left(x\right)$ $\widehat{b}_{t,L}\left(x\right)$
rMSE $\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left(\widehat{b}_{t,d}^{(m)}\left(x_{it}^{(m)}\right)-b_{t,d}\left(x_{it}^{(m)}\right)\right)^{2}}/\left|\overline{b}_{d}\right|$ 3.36% 5.57% 5.59%
MND $\left(NTM\right)^{-1}\sum_{i,t,m}\left|\widehat{b}_{t,d}^{(m)}\left(x_{it}^{(m)}\right)-b_{t,d}\left(x_{it}^{(m)}\right)\right|/\left|\overline{b}_{d}\right|$ 2.87% 4.36% 4.37%
To show how robust my method is in estimating $\overline{b}$ and $b_{t}\left(x\right)$, I conduct another set of exercises. I evaluate the performance of my TERC estimator using rMSE defined as
$\sqrt{M^{-1}\sum_{m}\left\|\widehat{\overline{b}}^{(m)}-\overline{b}\right\|^{2}}/\left\|\overline{b}\right\|$
for the vector $\overline{b}$, and
$\sqrt{\left(NTM\right)^{-1}\sum_{i,t,m}\left\|\widehat{b}_{t}^{(m)}\left(x_{it}^{(m)}\right)-b_{t}\left(x_{it}^{(m)}\right)\right\|^{2}}/\left\|\overline{b}\right\|$
@@
for the vector $b_{t}\left(x\right)$. First, I vary $n$ and $T$, and present the results in Table A.7. As expected, a larger $n$ is good for overall performance. However, the magnitude in the improvement of performance is mild, possibly because with more agents I need more data to control for the increasing degree of heterogeneity. On the other hand, I find that the method performs reasonably well even with a small sample size of $n=500$. Having a larger $T$ improves the performance, which is again expected as I can exploit more information from repeated observations of the same individual to better control for the fixed effect $A_{i}$.
Table A.7: Performance under Varying $n$ and $T$
@@
Second, I vary the degree of the spline bases used to construct the series estimators for all steps. The knot is placed at the median for all specifications. Table A.8 contains the results. I find that increasing the degree of basis functions from one to two improves estimation accuracy significantly. When I increase the degree of basis functions from two to three, the performance deteriorates. When the degree becomes larger, there are more regressors in each step of regression, leading to a possible over-fitting problem. Based on this result, I use the second-degree splines in the empirical application. Note that one may also use the AIC criterion to select the degree of basis functions.
Table A.8: Performance under Varying Degree of Basis Functions
@@
Lastly, I examine how including the ex-post shocks $\widetilde{\epsilon}_{it}$ and $\upsilon_{it}$ into the model affects the performance of my estimators. I draw $\widetilde{\epsilon}_{it}\sim\mathcal{U}\left[-.25,.25\right]$, the ex-post shock to the main equation (2.1), independently of all the other variables. $\widetilde{\epsilon}_{it}$ can also be interpreted as an ex-post shock to $\omega_{it}$. For $\upsilon_{it}$, since ex-post shock to $\omega_{it}$ has been considered by $\widetilde{\epsilon}_{it}$, I draw $\upsilon_{it}^{s}\sim\mathcal{U}\left[-.1,.1\right]$ independently of all the other variables and use $\beta_{it,s}^{\text{new}}=\beta_{it,s}+\upsilon_{it,s}$ in (2.1) for $s\in\left\{K,L\right\}$. Results are presented in Table A.9. As expected, adding ex-post errors $\widetilde{\epsilon}_{it}$ or $\upsilon_{it}$ negatively affects the performance of the estimators. When $\widetilde{\epsilon}_{it}$ is included, the rMSE of $\widehat{\overline{b}}$ for estimating $\overline{b}$ increases from 2.99% to 3.64% and the rMSE of $\widehat{b}_{t}\left(x\right)$ for estimating $b_{t}\left(x\right)$ rises from 3.64% to 5.43%. The effect of including $\upsilon_{it}$ on the performance is similar.
Table A.9: Performance with and without Ex Post Shocks
Add $\epsilon_{it}$ to $Y_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$ Add $\upsilon_{it}$ to $\beta_{it}$? rMSE: $\overline{b}$ rMSE: $b_{t}\left(x\right)$
No 2.99% 3.69% No 2.99% 3.69%
Yes 3.64% 5.43% Yes 3.57% 5.19%